A woman deposits consecutive equal half yearly amounts of R3400 into a savings account starting 6 months from now and ending 5 and a half years from now. The savings account offers an interest rate of 6.5%p.a. What is the future value of the savings account immediately after the last deposit was made?
1
Expert's answer
2020-05-25T20:54:27-0400

We solve by the formula

"FVA=A\times\frac{(1+\frac{i}{2})^{10}-1}{\frac{i}{2}}\times(1+\frac{i}{2})"

"FVA=3400\times\frac{(1+\frac{0.065}{2})^{10}-1}{\frac{0.065}{2}}\times(1+\frac{0.065}{2})"

FVA=40 710.38